OpenStudy (anonymous):

what is the probability of drawing P(Jack) in a standard set of cards? P(Spade)? P(Jack of spades)? P(not spade)?

OpenStudy (kirbykirby):

There are 4 suits (hearts, diamonds, spades, clubs) in a deck. And there is one Jack for every suit (so 4 Jacks in total) Since a normal deck has 52 cars, then P(Jack) = 4/52 ---- For P(Spade), notice that one quarter of the deck consists of spades... so P(spade) = \(\dfrac{52\div4}{52}\) ---- \(P(\text{Jack or Spade})=P(\text{Jack})+P(\text{Spade})-P(\text{Jack and Spade})\) (by the rule for unions) Now P(Jack and Spade) consists of the cards that are Both a Jack and a Spade. Can you figure out how many this is? ----- \(P(\text{not Spade)}=1-P(\text{Spade})\) (by the rule for complements)